OpenStudy (jazzyfa30):

help please b^3+8b^2-20b/b-2

OpenStudy (fifciol):

factor out b in numerator

OpenStudy (jazzyfa30):

so b^2+8b-20/2

OpenStudy (fifciol):

no it's\[\frac{ b(b^2+8b-20) }{b-2 }\]

OpenStudy (anonymous):

\[b^3+8b^2-20b/(b-2)?\]

OpenStudy (jazzyfa30):

ok

OpenStudy (anonymous):

...and you are solving for b?

OpenStudy (fifciol):

Remember to determine the domain- denominator must be different from 0

OpenStudy (jazzyfa30):

can u help me get the answer cus that didnt help any

OpenStudy (anonymous):

\[b 3 +8b 2 −20b/(b−2)\] b's cancel giving: \[=b^3+8b^2+10\]

OpenStudy (jazzyfa30):

ok what about the -2

OpenStudy (anonymous):

pull b^2 \[=b^2(b+8)+10\]

OpenStudy (anonymous):

the b's cancel and it leaves -20/-2. which = +10

OpenStudy (jazzyfa30):

ohhh

OpenStudy (anonymous):

Is that all you need help with? or do we need to go further?

OpenStudy (jazzyfa30):

so my answer is b2(b+8)+10

OpenStudy (anonymous):

yes, or the first answer I gave. they are both the same. just one is broken down further.

OpenStudy (jazzyfa30):

ok so can u help me with 4 more